# SRA-CL: Semantic Retrieval Augmented Contrastive Learning for Sequential Recommendation

### v4

@@ 1 Introduction
Sequential recommendation aims to model user preferences based on historical behavior sequences, a task of significant value for online platforms like YouTube and Amazon. However, accurate preference modeling faces a fundamental challenge: data sparsity, as most users have only limited interaction records and most items receive little attention. To address this issue, numerous self-supervised learning techniques [37, 44, 32] have been proposed, leveraging auxiliary tasks to improve data utilization efficiency. Among these, contrastive learning has emerged as a predominant approach due to its conceptual simplicity and proven effectiveness [32, 22, 3, 20, 45]. Typically, it constructs positive sample pairs from the data and maximizes their agreement in the representation space [2].
@@
As illustrated in Figure 1, existing contrastive learning approaches for sequential recommendation can be broadly classified into two categories: (1) inter-user contrastive learning, which contrasts sequences from different users, and (2) intra-user contrastive learning, which contrasts different augmented views of a single user’s sequence. In the inter-user paradigm, user sequence representations are clustered using K-means, and users within the same cluster are treated as positive samples for each other [3, 15, 21]. In the intra-user paradigm, perturbations are applied to a user’s sequence to generate augmented views, and the similarity between these views is maximized [32, 19, 22, 20]. These contrastive learning methods are typically employed as auxiliary tasks alongside the primary recommendation objective and have been demonstrated to enhance recommendation performance by improving user representation learning [22, 34].
@@
Despite their empirical success, existing methods suffer from several limitations in contrastive pair construction, which may undermine their effectiveness in recommendation scenarios. 1) Semantic Divergence. Many existing methods construct contrastive pairs through random augmentation operations such as random masking [32, 19] and Dropout [22]. However, in sequential recommendation where data is inherently sparse and exhibits sequential patterns, such random operations may lead to a complete change in the sequence’s semantics (i.e., user preferences). Bringing semantically different sequences closer together in embedding space may diminish the model’s ability to discriminate among distinct user preferences. Additionally, some methods determine contrastive pairs by clustering user representations derived from collaborative signals [3, 21], where users within the same cluster are considered positive pairs. However, the sparse ID signals can lead to low-quality representations and inaccurate clustering results. 2) Unlearnability. Existing methods rely on predefined rules to construct positive pairs, such as directly selecting users from the same cluster [3, 21, 19], or treating sequences sharing the same next item as positive pairs [22, 21]. These rigid heuristics impose strong assumptions that constrain models from autonomously learning optimal contrastive pairs. Moreover, the approach of using sequences with identical next items as positive pairs essentially replicates the recommendation objective (i.e., next-item prediction), providing no additional information gain. Therefore, the suboptimal construction of contrastive pairs in existing methods limits their effectiveness and hinders contrastive learning’s full potential.
@@
Given these limitations, constructing high-quality contrastive samples remains a critical challenge. Semantic information, which is readily available in textual data such as product categories, brands, and descriptions, provides a promising solution. Unlike sparse behavior signals, semantic data maintains validity regardless of data volume or training dynamics, as it derives from structured knowledge rather than co-occurrence patterns [44]. Additionally, semantic features offer complementary information beyond collaborative signals. Motivated by these advantages, we propose leveraging semantic information to construct superior contrastive pairs. However, accurately capturing user preferences requires models with powerful understanding and reasoning capabilities. Recent research has shown that large language models (LLMs) can effectively understand user preferences and achieve competitive performance on sequential recommendation tasks [33]. Inspired by this, we propose to enhance contrastive learning through LLM-powered semantic retrieval.
@@ 2.2 Sequential Recommendation Backbone
Our method is model-agnostic and can be integrated with various sequential recommendation models, as demonstrated in Section 4.3. To facilitate the introduction of our approach, we adopt the transformer architecture [26] as the backbone recommendation model following previous studies [20, 21, 22].
@@
  • Sequence Encoder. The representation of the input sequence is then fed into $L$ Transformer layers [26] to capture complex sequential patterns, which can be defined as follows:
@@
  • Prediction and Objective Function. During prediction, we calculate the probability of each item using $\mathbf{\hat{y}}=\text{softmax}(\mathbf{h}_{u}\mathbf{M}^{\mathrm{T}})$, where $\mathbf{\hat{y}}\in\mathbb{R}^{|\mathcal{V}|}$ and $\hat{y}_{v}$ represents the likelihood of item $v$ being the next item. For training, we adopt the same cross-entropy loss function as our baseline methods [20, 21, 22] to ensure fairness, where $v^{+}$ denotes the ground truth item for user $u$.
@@ 3.1 Inter-User Contrastive Learning via User Semantic Retrieval
    where $\tilde{\mathbf{h}}_{u}\in\mathbb{R}^{\tilde{d}}$ represents the semantic embedding of user preferences and $\tilde{d}$ is the embedding size of the text embedding model $\mathcal{M}$. Specifically, $\mathcal{M}$ indicates SimCSE-RoBERTa [8] in this paper due to its open-source availability and excellent sentence semantic extraction capabilities. The generated semantic embeddings are cached and remain fixed throughout the whole training process.
@@
      where $\mathbf{W}\in\mathbb{R}^{d\times\tilde{d}}$ is a learnable weight matrix, and $\|$ denotes the concatenation operation. $\mathbf{a}\in\mathbb{R}^{2d}$ represents a single-layer neural network used to generate the attention score, with the LeakyReLU activation function adopted [27]. The softmax function is employed to transform the coefficients into probabilities. Based on this, we generate the composite positive contrastive sample $\mathbf{h}_{u}^{+}$ for $\mathbf{h}_{u}$ by:
@@ 3.2 Intra-User Contrastive Learning via Item Semantic Retrieval
For intra-user contrastive learning, most existing methods apply predefined random perturbations to the original sequence to generate augmented views, which are treated as a pair of positive samples [32, 22]. A significant limitation of them is the introduction of considerable uncertainty in the semantic similarity between positive samples. This substantial variation in user sequence semantics among positive samples undermines the reliability of the contrastive learning process. To address this issue, we leverage a comprehensive understanding of both the semantic information of the item itself and the typical contexts in which the item appears. Based on this understanding, we replace certain items in the sequence with similar ones, resulting in semantic-consistent positive samples.
@@ 4.1 Experimental Settings
  • Datasets. Following previous studies [19, 32, 20], we conducted experiments on four public real-world datasets: Yelp, Amazon Sports, Beauty, and Office. The statistics for these datasets are presented in Table 3. More details about the datasets are shown in Appendix B.1.
  • Evaluation Metrics. To evaluate the performance of the models, we use widely recognized evaluation metrics: Hit Rate (HR) and Normalized Discounted Cumulative Gain (NDCG), follow previous studies  [39, 28, 9, 12]. The leave-one-out strategy is employed, where the last interaction is used for testing, the second-to-last interaction for validation, and the remaining interactions for training. To ensure an unbiased evaluation, we rank the prediction on the whole item set without sampling.
  • Baseline Methods. We compare our method with 13 baseline methods, categorized into three groups: 1) classical methods (GRU4Rec [10], SASRec [12], BERT4Rec [25]), 2) contrastive learning-based methods ($\rm{{S}^{3}\text{-}{Rec}}$ [44], CL4SRec [32], CoSeRec [19], ICLRec [3], DuoRec [22], MCLRec [20], ICSRec [21]), and 3) LLM-based methods (LRD [33], RLMRec [23], LLM-ESR [18]).
  • Implementation Details. All experiments are conducted with a single 32G V100 GPU. The embedding size is set to 64. We adopt the batch size of 256 and employ the Adam optimizer with a learning rate of 0.001. The dropout rate is set to 0.5 across all datasets. Following previous studies [33], we set the maximum sequence length to 20. The early stopping is applied if the metrics on the validation set do not improve over 10 consecutive epochs. For LLM, we use DeepSeek-V3 by invoking its API. We set the LLM’s temperature $\tau$ to 0 and top-$p$ to 0.001. For the text embedding model $\mathcal{M}$, we use the pre-trained RoBERTa from Hugging Face. Note that identical settings are adopted for our method and baselines that involve LLMs and text embeddings to ensure fairness. More implementation details can be found in Appendix B.3.
@@ 4.4 Ablation Study
Metric
w/o
CL
w/o
$\mathcal{L}_{CS}$
w/o
$\mathcal{L}_{IS}$
w/o
learn.
w/o
semantic
w/o
LLM
Ours
@@ 5 Related Work
  • Contrastive Learning in Sequential Recommendation. Contrastive learning has been successfully used to enhance sequential recommendation [38, 11, 36, 29, 31, 35, 4, 5, 21]. In terms of the composition of contrastive samples, we categorize existing methods into two types: (1) Inter-user. This involves generating contrastive samples from different user sequences. For example, ICLRec [3] clusters user interests into distinct categories by K-Means and brings the representations of users with similar interests closer. ICSRec [21] further segments a user’s behavior sequence into multiple subsequences to generate finer-grained user intentions for contrastive learning. These methods generate contrastive supervision signals based on collaborative signals. However, the sparsity of the co-occurrence pattern leads to unreliable clustering results, which in turn affects the performance of contrastive learning. (2) Intra-user. This involves applying perturbations to the original sequence to generate augmented views. The two views of the same sequence are treated as a pair of positive samples. For example, CL4SRec [32] employs three data-level augmentation operators: Cropping, Masking, and Reordering, to create contrastive pairs. CoSeRec [19] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec. In addition, some methods generate augmented views from the model’s hidden layers. A notable example is DuoRec [22], which creates positive pairs by forward-passing a sequence representation twice with different dropout masks. MCLRec [20] further combines data-level and model-level augmentation. Despite their effectiveness, they employ random operators, introducing significant uncertainty and potentially generating unreasonable positive samples for contrastive learning.
@@
  • Sequential Recommendation with LLMs. Recent studies have successfully integrated LLMs into recommender systems [30, 41, 14, 6]. Overall, LLMs are employed either as direct recommenders or as tools for extracting semantic information [13, 17, 40, 1, 24]. In the former approach, all inputs are converted into textual format, and the LLM generates recommendations based on its pre-trained knowledge or after undergoing supervised fine-tuning. Representative examples include LC-Rec [42], LLM-TRSR [43], and CALRec [16]. However, these methods rely on the inference process of large language models to generate recommendation results, which is computationally expensive and often challenging to deploy in practical scenarios. Another line of research [24, 28, 18, 33, 40, 1] leverages LLMs to process semantic information and incorporates it into traditional ID-based models. For example, SLIM [28] distills knowledge from large-scale LLMs into a smaller student LLM to improve the recommendation model. LLM-ESR [18] addresses the long-tail problem by leveraging collaborative signals and semantic information through dual-view modeling and self-distillation. LRD [33] utilizes the LLM to explore potential relations between items and reconstructs one item based on its relation to another. Unlike the aforementioned methods, our approach, grounded in the essence of contrastive learning, aims to construct more effective contrastive pairs with LLMs.
@@ B.1 Datasets
We conducted experiments on four public real-world datasets: Yelp, Sports, Beauty, and Office. The statistics for these datasets are presented in Table 3. These datasets cover a diverse range of application scenarios. The Yelp dataset [^1] https://www.yelp.com/dataset is widely used for business recommendations. The Sports, Beauty, and Office datasets are sourced from Amazon [^2] http://jmcauley.ucsd.edu/data/amazon/, one of the largest e-commerce platforms. Following previous studies [19, 32, 20], the users and items that have fewer than five interactions are removed.
@@ B.2 Baseline Methods
  • GRU4Rec [10] applies recurrent neural networks (RNN) to sequential recommendation.
  • SASRec [12] is the first work to utilize the self-attention mechanism for sequential recommendation.
  • BERT4Rec [25] employs the BERT [7] framework to capture the context information of user behaviors.
  • $\rm{\textbf{S}^{3}\text{-}\textbf{Rec}}$ [44] leverages four self-supervised objectives to uncover the inherent correlations within the data.
  • CL4SRec [32] proposes three random augmentation operators to generate positive samples for contrastive learning.
  • CoSeRec [19] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec.
  • ICLRec [3] clusters user interests into distinct categories and brings the representations of users with similar interests closer together.
  • DuoRec [22] combines a model-level dropout augmentation and a sampling strategy for choosing hard positive samples.
  • MCLRec [20] integrates CL4SRec’s random data augmentation for the input sequence and employs MLP layers for model-level augmentation.
  • ICSRec [21] is an improvement on ICLRec, further segmenting a user’s sequential behaviors into multiple subsequences to generate finer-grained user intentions for contrastive learning.
  • LRD [33] is an LLM-based method. It leverages LLMs to discover new relations between items and reconstructs one item based on the relation and another item.
  • RLMRec [23] utilizes LLMs to generate text profiles and combine their semantic embeddings with recommendation models.
  • LLM-ESR [18] is also an LLM-based method. It addresses the long-tail problem by simultaneously leveraging collaborative signals and semantic information through the dual-view modeling and self-distillation.

### v5

# Semantic Retrieval Augmented Contrastive Learning for Sequential Recommendation

@@ 1 Introduction
Sequential recommendation aims to model user preferences based on historical behavior sequences, a task of significant value for online platforms like YouTube and Amazon. However, accurate preference modeling faces a fundamental challenge: data sparsity, as most users have only limited interaction records and most items receive little attention. To address this issue, numerous self-supervised learning techniques [yu2023self, zhou2020s3, xie2022contrastive] have been proposed, leveraging auxiliary tasks to improve data utilization efficiency. Among these, contrastive learning has emerged as a predominant approach due to its conceptual simplicity and proven effectiveness [xie2022contrastive, qiu2022contrastive, chen2022intent, qin2023meta, zhou2023equivariant]. Typically, it constructs positive sample pairs from the data and maximizes their agreement in the representation space [chen2020simple].
@@
As illustrated in Figure 1, existing contrastive learning approaches for sequential recommendation can be broadly classified into two categories: (1) inter-user contrastive learning, which contrasts sequences from different users, and (2) intra-user contrastive learning, which contrasts different augmented views of a single user’s sequence. In the inter-user paradigm, user sequence representations are clustered using K-means, and users within the same cluster are treated as positive samples for each other [chen2022intent, li2023multi, qin2024intent]. In the intra-user paradigm, perturbations are applied to a user’s sequence to generate augmented views, and the similarity between these views is maximized [xie2022contrastive, liu2021contrastive, qiu2022contrastive, qin2023meta]. These contrastive learning methods are typically employed as auxiliary tasks alongside the primary recommendation objective and have been demonstrated to enhance recommendation performance by improving user representation learning [qiu2022contrastive, yang2023debiased].
@@
Despite their empirical success, existing methods suffer from several limitations in contrastive pair construction, which may undermine their effectiveness in recommendation scenarios. 1) Semantic Divergence. Many existing methods construct contrastive pairs through random augmentation operations such as random masking [xie2022contrastive, liu2021contrastive] and Dropout [qiu2022contrastive]. However, in sequential recommendation where data is inherently sparse and exhibits sequential patterns, such random operations may lead to a complete change in the sequence’s semantics (i.e., user preferences). Bringing semantically different sequences closer together in embedding space may diminish the model’s ability to discriminate among distinct user preferences. Additionally, some methods determine contrastive pairs by clustering user representations derived from collaborative signals [chen2022intent, qin2024intent], where users within the same cluster are considered positive pairs. However, the sparse ID signals can lead to low-quality representations and inaccurate clustering results. 2) Unlearnability. Existing methods rely on predefined rules to construct positive pairs, such as directly selecting users from the same cluster [chen2022intent, qin2024intent, liu2021contrastive], or treating sequences sharing the same next item as positive pairs [qiu2022contrastive, qin2024intent]. These rigid heuristics impose strong assumptions that constrain models from autonomously learning optimal contrastive pairs. Moreover, the approach of using sequences with identical next items as positive pairs essentially replicates the recommendation objective (i.e., next-item prediction), providing no additional information gain. Therefore, the suboptimal construction of contrastive pairs in existing methods limits their effectiveness and hinders contrastive learning’s full potential.
@@
Given these limitations, constructing high-quality contrastive samples remains a critical challenge. Semantic information, which is readily available in textual data such as product categories, brands, and descriptions, provides a promising solution. Unlike sparse behavior signals, semantic data maintains validity regardless of data volume or training dynamics, as it derives from structured knowledge rather than co-occurrence patterns [zhou2020s3]. Additionally, semantic features offer complementary information beyond collaborative signals. Motivated by these advantages, we propose leveraging semantic information to construct superior contrastive pairs. However, accurately capturing user preferences requires models with powerful understanding and reasoning capabilities. Recent research has shown that large language models (LLMs) can effectively understand user preferences and achieve competitive performance on sequential recommendation tasks [yang2024sequential]. Inspired by this, we propose to enhance contrastive learning through LLM-powered semantic retrieval.
@@ 2.2 Sequential Recommendation Backbone
Our method is model-agnostic and can be integrated with various sequential recommendation models, as demonstrated in Section 4.3. To facilitate the introduction of our approach, we adopt the transformer architecture [vaswani2017attention] as the backbone recommendation model following previous studies [qin2023meta, qin2024intent, qiu2022contrastive].
@@
  • Sequence Encoder. The representation of the input sequence is then fed into $L$ Transformer layers [vaswani2017attention] to capture complex sequential patterns, which can be defined as follows:
@@
  • Prediction and Objective Function. During prediction, we calculate the probability of each item using $\mathbf{\hat{y}}=\text{softmax}(\mathbf{h}_{u}\mathbf{M}^{\mathrm{T}})$, where $\mathbf{\hat{y}}\in\mathbb{R}^{|\mathcal{V}|}$ and $\hat{y}_{v}$ represents the likelihood of item $v$ being the next item. For training, we adopt the same cross-entropy loss function as our baseline methods [qin2023meta, qin2024intent, qiu2022contrastive] to ensure fairness, where $v^{+}$ denotes the ground truth item for user $u$.
@@ 3.1 Inter-User Contrastive Learning via User Semantic Retrieval
    where $\tilde{\mathbf{h}}_{u}\in\mathbb{R}^{\tilde{d}}$ represents the semantic embedding of user preferences and $\tilde{d}$ is the embedding size of the text embedding model $\mathcal{M}$. Specifically, $\mathcal{M}$ indicates SimCSE-RoBERTa [gao2021simcse] in this paper due to its open-source availability and excellent sentence semantic extraction capabilities. The generated semantic embeddings are cached and remain fixed throughout the whole training process.
@@
      where $\mathbf{W}\in\mathbb{R}^{d\times\tilde{d}}$ is a learnable weight matrix, and $\|$ denotes the concatenation operation. $\mathbf{a}\in\mathbb{R}^{2d}$ represents a single-layer neural network used to generate the attention score, with the LeakyReLU activation function adopted [velivckovic2017graph]. The softmax function is employed to transform the coefficients into probabilities. Based on this, we generate the composite positive contrastive sample $\mathbf{h}_{u}^{+}$ for $\mathbf{h}_{u}$ by:
@@ 3.2 Intra-User Contrastive Learning via Item Semantic Retrieval
For intra-user contrastive learning, most existing methods apply predefined random perturbations to the original sequence to generate augmented views, which are treated as a pair of positive samples [xie2022contrastive, qiu2022contrastive]. A significant limitation of them is the introduction of considerable uncertainty in the semantic similarity between positive samples. This substantial variation in user sequence semantics among positive samples undermines the reliability of the contrastive learning process. To address this issue, we leverage a comprehensive understanding of both the semantic information of the item itself and the typical contexts in which the item appears. Based on this understanding, we replace certain items in the sequence with similar ones, resulting in semantic-consistent positive samples.
@@ 4.1 Experimental Settings
  • Datasets. Following previous studies [liu2021contrastive, xie2022contrastive, qin2023meta], we conducted experiments on four public real-world datasets: Yelp, Amazon Sports, Beauty, and Office. The statistics for these datasets are presented in Table 3. More details about the datasets are shown in Appendix B.1.
  • Evaluation Metrics. To evaluate the performance of the models, we use widely recognized evaluation metrics: Hit Rate (HR) and Normalized Discounted Cumulative Gain (NDCG), follow previous studies  [zhang2024finerec, wang2024can, he2020lightgcn, kang2018self]. The leave-one-out strategy is employed, where the last interaction is used for testing, the second-to-last interaction for validation, and the remaining interactions for training. To ensure an unbiased evaluation, we rank the prediction on the whole item set without sampling.
  • Baseline Methods. We compare our method with 13 baseline methods, categorized into three groups: 1) classical methods (GRU4Rec [hidasi2015session], SASRec [kang2018self], BERT4Rec [sun2019bert4rec]), 2) contrastive learning-based methods ($\rm{{S}^{3}\text{-}{Rec}}$ [zhou2020s3], CL4SRec [xie2022contrastive], CoSeRec [liu2021contrastive], ICLRec [chen2022intent], DuoRec [qiu2022contrastive], MCLRec [qin2023meta], ICSRec [qin2024intent]), and 3) LLM-based methods (LRD [yang2024sequential], RLMRec [ren2024representation], LLM-ESR [liu2024llm]).
  • Implementation Details. All experiments are conducted with a single 32G V100 GPU. The embedding size is set to 64. We adopt the batch size of 256 and employ the Adam optimizer with a learning rate of 0.001. The dropout rate is set to 0.5 across all datasets. Following previous studies [yang2024sequential], we set the maximum sequence length to 20. The early stopping is applied if the metrics on the validation set do not improve over 10 consecutive epochs. For LLM, we use DeepSeek-V3 by invoking its API. We set the LLM’s temperature $\tau$ to 0 and top-$p$ to 0.001. For the text embedding model $\mathcal{M}$, we use the pre-trained RoBERTa from Hugging Face. Note that identical settings are adopted for our method and baselines that involve LLMs and text embeddings to ensure fairness. More implementation details can be found in Appendix B.3.
@@ 4.4 Ablation Study
Metric w/o CL w/o $\mathcal{L}_{CS}$ w/o $\mathcal{L}_{IS}$ w/o learn. w/o semantic w/o LLM Ours
@@ 5 Related Work
  • Contrastive Learning in Sequential Recommendation. Contrastive learning has been successfully used to enhance sequential recommendation [zhang2024recdcl, jiang2023diffkg, yu2022graph, wu2021self, xia2023automated, yin2022autogcl, cui2022dual, cui2024context, qin2024intent]. In terms of the composition of contrastive samples, we categorize existing methods into two types: (1) Inter-user. This involves generating contrastive samples from different user sequences. For example, ICLRec [chen2022intent] clusters user interests into distinct categories by K-Means and brings the representations of users with similar interests closer. ICSRec [qin2024intent] further segments a user’s behavior sequence into multiple subsequences to generate finer-grained user intentions for contrastive learning. These methods generate contrastive supervision signals based on collaborative signals. However, the sparsity of the co-occurrence pattern leads to unreliable clustering results, which in turn affects the performance of contrastive learning. (2) Intra-user. This involves applying perturbations to the original sequence to generate augmented views. The two views of the same sequence are treated as a pair of positive samples. For example, CL4SRec [xie2022contrastive] employs three data-level augmentation operators: Cropping, Masking, and Reordering, to create contrastive pairs. CoSeRec [liu2021contrastive] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec. In addition, some methods generate augmented views from the model’s hidden layers. A notable example is DuoRec [qiu2022contrastive], which creates positive pairs by forward-passing a sequence representation twice with different dropout masks. MCLRec [qin2023meta] further combines data-level and model-level augmentation. Despite their effectiveness, they employ random operators, introducing significant uncertainty and potentially generating unreasonable positive samples for contrastive learning.
@@
  • Sequential Recommendation with LLMs. Building upon foundations laid by traditional recommender systems [he2020lightgcn, li2023alpt, li2024optfp], recent studies have successfully integrated LLMs into the recommender paradigm [wu2024survey, zhao2023recommender, li2024large, cui2025comprehending]. Overall, LLMs are employed either as direct recommenders or as tools for extracting semantic information [li2023prompt, liao2024llara, zhao2024let, boz2024improving, ren2024enhancing]. In the former approach, all inputs are converted into textual format, and the LLM generates recommendations based on its pre-trained knowledge or after undergoing supervised fine-tuning. Representative examples include LC-Rec [zheng2024adapting], LLM-TRSR [zheng2024harnessing], and CALRec [li2024calrec]. However, these methods rely on the inference process of large language models to generate recommendation results, which is computationally expensive and often challenging to deploy in practical scenarios. Another line of research [ren2024enhancing, wang2024can, liu2024llm, yang2024sequential, zhao2024let, boz2024improving] leverages LLMs to process semantic information and incorporates it into traditional ID-based models. For example, SLIM [wang2024can] distills knowledge from large-scale LLMs into a smaller student LLM to improve the recommendation model. LLM-ESR [liu2024llm] addresses the long-tail problem by leveraging collaborative signals and semantic information through dual-view modeling and self-distillation. LRD [yang2024sequential] utilizes the LLM to explore potential relations between items and reconstructs one item based on its relation to another. Unlike the aforementioned methods, our approach, grounded in the essence of contrastive learning, aims to construct more effective contrastive pairs with LLMs.
@@ B.1 Datasets
We conducted experiments on four public real-world datasets: Yelp, Sports, Beauty, and Office. The statistics for these datasets are presented in Table 3. These datasets cover a diverse range of application scenarios. The Yelp dataset [^1] https://www.yelp.com/dataset is widely used for business recommendations. The Sports, Beauty, and Office datasets are sourced from Amazon [^2] http://jmcauley.ucsd.edu/data/amazon/, one of the largest e-commerce platforms. Following previous studies [liu2021contrastive, xie2022contrastive, qin2023meta], the users and items that have fewer than five interactions are removed.
@@ B.2 Baseline Methods
  • GRU4Rec [hidasi2015session] applies recurrent neural networks (RNN) to sequential recommendation.
  • SASRec [kang2018self] is the first work to utilize the self-attention mechanism for sequential recommendation.
  • BERT4Rec [sun2019bert4rec] employs the BERT [devlin2018bert] framework to capture the context information of user behaviors.
  • $\rm{\textbf{S}^{3}\text{-}\textbf{Rec}}$ [zhou2020s3] leverages four self-supervised objectives to uncover the inherent correlations within the data.
  • CL4SRec [xie2022contrastive] proposes three random augmentation operators to generate positive samples for contrastive learning.
  • CoSeRec [liu2021contrastive] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec.
  • ICLRec [chen2022intent] clusters user interests into distinct categories and brings the representations of users with similar interests closer together.
  • DuoRec [qiu2022contrastive] combines a model-level dropout augmentation and a sampling strategy for choosing hard positive samples.
  • MCLRec [qin2023meta] integrates CL4SRec’s random data augmentation for the input sequence and employs MLP layers for model-level augmentation.
  • ICSRec [qin2024intent] is an improvement on ICLRec, further segmenting a user’s sequential behaviors into multiple subsequences to generate finer-grained user intentions for contrastive learning.
  • LRD [yang2024sequential] is an LLM-based method. It leverages LLMs to discover new relations between items and reconstructs one item based on the relation and another item.
  • RLMRec [ren2024representation] utilizes LLMs to generate text profiles and combine their semantic embeddings with recommendation models.
  • LLM-ESR [liu2024llm] is also an LLM-based method. It addresses the long-tail problem by simultaneously leveraging collaborative signals and semantic information through the dual-view modeling and self-distillation.